OpenStudy (mari1234):

Line WX contains (−1, 2) and (4, 12) Line YZ contains points (5, 8) and (−2, −6). Lines WX and YZ are perpendicular, because the slopes are the same parallel, because the product of the slopes is −1 perpendicular, because the product of the slopes is −1 parallel, because the slopes are the same

RhondaSommer (rhondasommer):

I can help

RhondaSommer (rhondasommer):

do you know how to get slope

OpenStudy (mari1234):

no i forgot

RhondaSommer (rhondasommer):

#WelcomeToOpenStudy :) My name is Rhonda and I will medal and fan you :)

OpenStudy (mari1234):

thank you

RhondaSommer (rhondasommer):

okay so \[-1 = x^1\]\[2=y^1\]\[4=x^2\]\[ 12=y^2\] for WX line slope= \[y^2-y^1/x^2-x^1\]

OpenStudy (mari1234):

i dont get it XD

OpenStudy (mari1234):

nvm i got it

RhondaSommer (rhondasommer):

okay sorry i had to go super quick XD